OpenStudy (anonymous):

an equation of the line tangent to the graph of f(x)=x(1-2x)^3 at the point (1,-1) is ..

OpenStudy (adamaero):

Do you know what a tangent line is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just need some help in finding the equation

OpenStudy (zale101):

Take the derivative first.

OpenStudy (zale101):

Do you know how? :)

OpenStudy (anonymous):

lol uhm sort of.. hehe. uhm so far i've got 3x(1-2x)^2(-2)

OpenStudy (zale101):

\[ f(x)=x(1-2x)^3\] \[\frac{ dy }{ dx }=\frac{ d }{ dx }[x(1-2x)^3]\] Use the product rule \(x*(1-2x)^3\) hence x is being multiplied by (1-2x)^3, and then use the chain rule when you take the derivative of (1-2x)^3. \[\frac{ dy }{ dx }=1*(1-2x)^3 + 3x(1-2x)^{2}(-2)]\] \[\frac{ dy }{ dx }=1*(1-2x)^3 -6x(1-2x)^{2}]\]

OpenStudy (zale101):

now, the equation of the tangent line is already given, to find the slope you plug in the given x from (1,-1) which is 1 f'(1)=? will be the slope then use the point slope formula \(y-y_1=m(x-x_1)\) where m is the slope you have solved for when f'(1)=?

OpenStudy (anonymous):

wait but why did u change the x into a 1. Isn't the product rule uv'+vu' so the first multiplied by the derivative of the 2nd?

OpenStudy (zale101):

because the derivative of x is 1 \[\frac{ d }{ dx }[1x^1]=1*1x^0=1\] member the power rule?

OpenStudy (zale101):

power rule is one of the short cuts for taking the derivative, where you have to multiply the exponent by the coefficient and subtract one from the exponent \[\frac{ dy }{ dx }=nx^{n-1}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

wait so i plug in one into the original equation?

OpenStudy (zale101):

One thing i forgot the mention, after you solvd for he slope, you plug the slope to m and the given point (1,-1) to (x1,y1) in the point slope formula

OpenStudy (zale101):

plug 1 to the derivative to get the slope

OpenStudy (anonymous):

okay so I got 7 for the slope

OpenStudy (anonymous):

plug 2 to the derivative to get the slope